[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jun 22, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jun 22 17:25:59 UTC 2017


HSU Shift Log for Thu Jun 22 2017 13:25:53 GMT-0400 (EDT)

Shift: H

Forecaster: L Avila

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues: A ticket has to be open related to the national TCV.  
There are some technical issues when warnings were dropped in the last  
advisory. Shannon and DZ are in the loop.

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170622/ff2fcac6/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list